OpenStudy (anonymous):

Simplify .. x^-4y^3x^4y^2

OpenStudy (anonymous):

\[x^{-4}\times y^{3}\times x^{4}\times y^{2}\]when multiplying same variables, add the exponents: rearrange: \(x^{-4}\times x^4\times y ^3\times y^2\) add exponents: \(x^0\times y^5\) simplify: \(y^5\)